# OAT-FM: Optimal Acceleration Transport for Improved Flow Matching

### v1

@@ 4 Experiments
4.1 Testing on the Low-dimensional OT Benchmark
We first validate OAT-FM on the low-dimensional OT benchmark (Tong et al., 2024), which includes five 2D point cloud transport tasks. After applying FM (Lipman et al., 2023), I/SB/OT-CFM (Tong et al., 2024), and VP-CFM (Albergo et al., 2023), respectively, to train an MLP-based generator by 20,000 batches, we leverage OAT-FM to continually refine the model with the same number of batches. For fairness, we compare the models achieved by the two-phase FM paradigm with those trained by the baseline methods (i.e., FM, I/SB/OT-CFM, and VP-CFM) with 40,000 batches on two metrics: $i)$ the 2-Wasserstein distance $\mathcal{W}_{2}^{2}(\hat{\rho}_{1},\rho_{1})$ between the distribution of generated samples $\hat{\rho}_{1}$ and that of target data $\rho_{1}$, and ii)ii) the Normalized Path Energy (NPE) defined in terms of the 2-Wasserstein distance as
Here, we generate all samples using the RK45 ODE solver (Dormand and Prince, 1980) with 101 integration steps from $t=0$ to $t=1$ and compute $\mathrm{PE}$ accordingly. This metric evaluates the transport cost of the learned flow relative to the dynamic optimal transport.
@@ 4.1 Testing on the Low-dimensional OT Benchmark
Figure 2: An illustration of refining the flow of I-CFM via OAT-FM on the eight Gaussians to the Moons dataset. We conduct this experiment based on the code base provided in (Tong et al., 2024).
@@ 4.2 Unconditional Image Generation
Table 3: Comparisons of various methods in unconditional CIFAR-10 image generation. In the column “#Batch”, the number of training batches of each baseline method is in black, while that of OAT-FM is in purple. The unit “K” means 1,000 batches. The results of the methods labeled by $\ast$” are from Lee et al. (2024).
@@
DDPM (Ho et al., 2020) 1,000 3.17
Score SDE (Song et al., 2021b) 2,000 2.38
LSGM (Vahdat et al., 2021) 147 2.10
2-ReFlow++ (Lee et al., 2024) 35 2.30
EDM (Karras et al., 2022) 35 1.96
@@
Beyond the above low-dimensional OT benchmark, we further evaluate OAT-FM on generating CIFAR-10 images (Krizhevsky, 2009). We initialize our training from the pre-trained FM (Lipman et al., 2023), I-CFM (Tong et al., 2024), OT-CFM (Tong et al., 2024) and EDM (Karras et al., 2022) models, which serve as a starting point by providing good estimates for the boundary velocities. For FM, I-CFM, and OT-CFM, we follow the settings in (Tong et al., 2024). For EDM, same as (Lee et al., 2024), we adapt it into a flow matching model by adjusting its time and scaling factors, which allows a seamless transition to our phase-2 training (see Appendix C.1). All models are trained with a per-GPU batch size of 128 and an EMA decay rate of 0.9999. In the inference phase, we follow the standard setting, employing the Dopri5 solver for FM and I/OT-CFM and the Heun solver for EDM, respectively. For each model, we evaluate image quality using the Fréchet Inception Distance (FID) (Heusel et al., 2017) and record its number of training batches and that of inference steps (denoted as NFE) as well.
As shown in Table 3, OAT-FM consistently enhances the generation quality across all pre-trained models. Notably, for FM, I-CFM, and OT-CFM, our method achieves superior FID scores while requiring only 1K additional training batches, a substantial reduction from the 400K batches used by the original models. Furthermore, OAT-FM also improves upon the strong EDM baseline, lowering the FID from 1.96 to 1.93 with only 12K additional training batches. This result is also better than other competitive generative modeling methods, including the strong two-phase FM method 2-ReFlow++ (Lee et al., 2024). These results underscore that OAT-FM serves as an effective and computationally efficient plug-in module for refining existing unconditional generative models, thereby boosting their performance with minimal training overhead.
@@
Lower-level Upper-level Phase-1 Method
Problem Problem FM EDM
Without Phase-2 Training 3.71 1.96
$\mathcal{W}_{2}^{2}$ in (3) $\mathcal{L}_{\text{CFM}}$ in (3) 3.75 8.77
$\mathcal{W}_{2}^{2}$ in (3) $\mathcal{L}_{\text{OAT}}$ in (9) 3.55 8.68
$\mathcal{A}_{2}^{2}$ in (9) $\mathcal{L}_{\text{CFM}}$ in (3) 3.81 1.95
$\mathcal{A}_{2}^{2}$ in (9) $\mathcal{L}_{\text{OAT}}$ in (9) 3.54 1.93
Ablation Studies. As shown in Table 4, we conduct ablation studies on CIFAR-10 to validate the design of OAT-FM. In particular, given the models trained by FM (Lipman et al., 2023) and EDM (Karras et al., 2022), respectively, we continually train them by OAT-FM under different settings, dissecting the contributions of its key components, i.e., OAT-based coupling computation in its lower-level problem and the OAT-based objective used in its upper-level problem. When deriving the coupling by computing the Wasserstein distance in (3), the FM-based model maintains its low FID score on CIFAR-10 while EDM suffers severe performance degradation no matter what upper-level objective is. For the upper-level objective, we can find that $\mathcal{L}_{\text{OAT}}$ works better than $\mathcal{L}_{\text{CFM}}$ consistently for both FM- and EDM-based models. These results demonstrate that each component of OAT-FM plays an indispensable role in effectively refining flows and enhancing model performance.
@@ 4.3 Large-scale Conditional Image Generation
To verify the feasibility of OAT-FM as a phase-2 training method in practice, given the state-of-the-art image generator SiT-XL (Ma et al., 2024) trained on the ImageNet $\times256$ benchmark (Deng et al., 2009), we apply OAT-FM to continually refine the model on the dataset and test it in conditional image generation tasks. We compare the model refined by OAT-FM with the original SiT-XL and other image generators on different metrics, including FID, spatial FID (sFID), Inception Score (IS) (Salimans et al., 2016), and the precision and recall measuring how well the real and generated data manifolds are overlapped with each other. In the inference phase, we apply an ODE sampler and a SDE one, respectively, to generate images with different Classifier-Free Guidance (CFG) scales.
As detailed in Table 5, for the original SiT-XL model derived by 1,400 training epochs, refining it by OAT-FM with merely five more training epochs (48K batches) leads to consistent improvements in FID, sFID, and IS. The precision and recall remain stable. In Figure 3, we plot the performance of SiT-XL with and without OAT-FM while varying the CFG scale from 1.0 (no guidance) to 4.0. The results confirm that OAT-FM delivers consistent improvements across the entire range of scales. Note that, setting a large CFG scale helps generate high-quality images in general (Ma et al., 2024; Peebles and Xie, 2023) (although leading to relatively high FID/sFID scores). Based on this fact, the superiority of OAT-FM is significant when CFG is large (e.g., CFG is 4.0), which further demonstrates the effectiveness of OAT-FM in practice.
@@ 5 Conclusion
BigGAN-deep (Brock et al., 2019) 6.95 7.36 171.4 0.87 0.28
StyleGAN-XL (Sauer et al., 2022) 2.30 4.02 265.1 0.78 0.53
Mask-GIT (Chang et al., 2022) 6.18 - 182.1 - -
ADM-G/U (Dhariwal and Nichol, 2021) 3.94 6.14 215.8 0.83 0.53
CDM (Ho et al., 2022) 4.88 - 158.7 - -
RIN (Jabri et al., 2023) 3.42 - 182.0 - -
@@
(a) FID and sFID
(b) Image Quality
Figure 3: The performance of SiT-XL + OAT-FM with different classifier-free guidance scales. (a) shows the comparison on FID and sFID scores. (b) shows the generated images when CFG is 4.0.
Limitations and future work. There are natural directions to refine the method. OAT-FM benefits from reasonably accurate endpoint velocities, so Currently, training from scratch via OAT-FM can be fragile due to its dependence on velocity information — early velocity estimates are noisy and can misguide the product-space coupling. As a straightforward solution, warm starts using CFM or self-distillation in the spirit of Shortcut (Frans et al., 2025) and consistency models (Kim et al., 2024) may offer simple remedies before handing off to OAT-FM. However, whether the theory of OAT can be used to develop one-step FM method is still an open problem. From a computational standpoint, mini-batch couplings scale quadratically in batch size. In the future, we plan to explore the dual form of OAT problem and develop a more efficient OAT solver to accelerate OAT-FM further.
@@ B.1 Derivation of (9)
In the context of FM, we have $\pi(z_{0},z_{1})=\pi(x_{0},v_{0},x_{1},v_{1})=\pi_{x}(x_{0},x_{1})\pi(v_{0}|x_{0})\pi(v_{1}|x_{1})$. Here, the sample-level coupling $\pi_{x}(x_{0},x_{1})\in\Pi(\rho_{0},\rho_{1})$ is the marginal coupling corresponding to sample pairs and $\pi(\cdot|x_{t})=\delta_{v_{\theta}(x_{t},t)}$ is the Dirac measure determined by the flow model $v_{\theta}$. As a result, we can merely optimize $\pi_{x}(x_{0},x_{1})$, and the lower-level problem in (7) becomes
@@
where $\langle\cdot,\cdot\rangle$ denotes inner product, $\mathbf{T}$ is the coupling matrix, and $\mathbf{C}=[c_{ij}]\in\mathbb{R}^{B\times B}$ is the cost matrix, whose element $c_{ij}=\|x_{1,i}-x_{0,j}\|_{2}^{2}-(x_{1,i}-x_{0,j})^{\top}(v_{0,j}+v_{1,i})$. Following the OT-CFM rationale, the problem can be cast as a linear program with computational complexity $\mathcal{O}(B^{3}\log\|\mathbf{C}\|_{\infty})$. The optimizer can be further approximated by adding an entropic regularizer of $\mathbf{T}$ weighted by $\epsilon$, i.e., $\epsilon\langle\mathbf{T},\log\mathbf{T}\rangle$. This allows the problem to be solved efficiently by Sinkhorn method, whose complexity is $\mathcal{O}(B^{2}\log B)$, and the exact OT result can be recovered by taking $\epsilon$ sufficiently small. We refer to Peyré and Cuturi (2019) for further details.
@@ Appendix C Detailed Experiment Settings
Model Paradigm Objective Path / Coupling Backbone Space
EDM (Karras et al., 2022) Diffusion Score Matching VP/VE Noise Schedule U-Net Pixel
FM (Lipman et al., 2023) Flow Flow Matching Gaussian-Data Path U-Net Pixel
I-CFM (Tong et al., 2024) Flow Conditional FM Linear Path (Independent) U-Net Pixel
VP-CFM (Albergo et al., 2023) Flow Conditional FM Trigonometric Path U-Net Pixel
SB-CFM (Tong et al., 2024) Flow Conditional FM Brownian Bridge Path U-Net Pixel
OT-CFM (Tong et al., 2024) Flow Conditional FM Linear Path (OT Coupling) U-Net Pixel
SiT-XL (Ma et al., 2024) Flow Velocity Matching Linear / GVP Path DiT Latent
@@ C.1 Conversion from VP/VE Models to Flow Matching
Following the work in (Lee et al., 2024), our training process begins with a pre-trained diffusion model. Specifically, we use models trained with the Elucidating the Design Space of Diffusion-Based Generative Models (EDM) framework (Karras et al., 2022), which provides a unified perspective on score-based models including Variance Preserving (VP) and Variance Exploding (VE) SDEs. To adapt this model for flow matching, we re-parameterizes its input and output to function as a velocity field predictor. This allows us to initialize our model with a strong, pre-trained diffusion backbone, facilitating a seamless transition to the second-phase OAT-FM training paradigm.
EDM Preconditioning. The EDM model takes a noise-corrupted input $x_{\sigma}=x_{1}+\sigma z$ (where $z\sim\mathcal{N}(0,I)$) and a noise level $\sigma$, is trained to denoise it. The denoised output $D_{\theta}(x_{\sigma},\sigma)$ is an estimate of the original data $x_{1}$ and is formulated using a preconditioning scheme:

### v2

@@ 4 Experiments
4.1 Testing on the Low-dimensional OT Benchmark
We first validate OAT-FM on the low-dimensional OT benchmark (tong2024improving), which includes five 2D point cloud transport tasks. After applying FM (lipman2023flow), I/SB/OT-CFM (tong2024improving), and VP-CFM (albergo2023stochastic), respectively, to train an MLP-based generator by 20,000 batches, we leverage OAT-FM to continually refine the model with the same number of batches. For fairness, we compare the models achieved by the two-phase FM paradigm with those trained by the baseline methods (i.e., FM, I/SB/OT-CFM, and VP-CFM) with 40,000 batches on two metrics: $i)$ the 2-Wasserstein distance $\mathcal{W}_{2}^{2}(\hat{\rho}_{1},\rho_{1})$ between the distribution of generated samples $\hat{\rho}_{1}$ and that of target data $\rho_{1}$, and ii)ii) the Normalized Path Energy (NPE) defined in terms of the 2-Wasserstein distance as
Here, we generate all samples using the RK45 ODE solver (dormand1980family) with 101 integration steps from $t=0$ to $t=1$ and compute $\mathrm{PE}$ accordingly. This metric evaluates the transport cost of the learned flow relative to the dynamic optimal transport.
Figure 2: An illustration of refining the flow of I-CFM via OAT-FM on the eight Gaussians to the Moons dataset. We conduct this experiment based on the code base provided in (tong2024improving).
@@ 4.2 Unconditional Image Generation
Beyond the above low-dimensional OT benchmark, we further evaluate OAT-FM on generating CIFAR-10 images (krizhevsky2009learning). We initialize our training from the pre-trained FM (lipman2023flow), I-CFM (tong2024improving), OT-CFM (tong2024improving) and EDM (karras2022elucidating) models, which serve as a starting point by providing good estimates for the boundary velocities. For FM, I-CFM, and OT-CFM, we follow the settings in (tong2024improving). For EDM, same as (lee2024improving), we adapt it into a flow matching model by adjusting its time and scaling factors, which allows a seamless transition to our phase-2 training (see Appendix C.1). All models are trained with a per-GPU batch size of 128 and an EMA decay rate of 0.9999. In the inference phase, we follow the standard setting, employing the Dopri5 solver for FM and I/OT-CFM and the Heun solver for EDM, respectively. For each model, we evaluate image quality using the Fréchet Inception Distance (FID) (heusel2017gans) and record its number of training batches and that of inference steps (denoted as NFE) as well.
As shown in Table 3, OAT-FM consistently enhances the generation quality across all pre-trained models. Notably, for FM, I-CFM, and OT-CFM, our method achieves superior FID scores while requiring only 1K additional training batches, a substantial reduction from the 400K batches used by the original models. Furthermore, OAT-FM also improves upon the strong EDM baseline, lowering the FID from 1.96 to 1.93 with only 12K additional training batches. This result is also better than other competitive generative modeling methods, including the strong two-phase FM method 2-ReFlow++ (lee2024improving). These results underscore that OAT-FM serves as an effective and computationally efficient plug-in module for refining existing unconditional generative models, thereby boosting their performance with minimal training overhead.
Table 3: Comparisons of various methods in unconditional CIFAR-10 image generation. In the column “#Batch”, the number of training batches of each baseline method is in black, while that of OAT-FM is in purple. The unit “K” means 1,000 batches. The results of the methods labeled by $\ast$” are from lee2024improving.
@@
DDPM 1,000 3.17
Score SDE 2,000 2.38
LSGM 147 2.10
2-ReFlow++ 35 2.30
EDM 35 1.96
@@
Lower Upper Phase-1
Prob. Prob. FM EDM
W/O Phase-2 3.71 1.96
$\mathcal{W}_{2}^{2}$ $\mathcal{L}_{\text{CFM}}$ 3.75 8.77
$\mathcal{W}_{2}^{2}$ $\mathcal{L}_{\text{OAT}}$ 3.55 8.68
$\mathcal{A}_{2}^{2}$ $\mathcal{L}_{\text{CFM}}$ 3.81 1.95
$\mathcal{A}_{2}^{2}$ $\mathcal{L}_{\text{OAT}}$ 3.54 1.93
Figure 3: The comparison of SiT-XL and SiT-XL + OAT-FM on FID and sFID.
Ablation Studies. As shown in Table 4, we conduct ablation studies on CIFAR-10 to validate the design of OAT-FM. In particular, given the models trained by FM (lipman2023flow) and EDM (karras2022elucidating), respectively, we continually train them by OAT-FM under different settings, dissecting the contributions of its key components, i.e., OAT-based coupling computation in its lower-level problem and the OAT-based objective used in its upper-level problem. To adapt EDM for flow matching, we reparameterize its denoising as a velocity field predictor. This allows us to initialize our model with a strong, pre-trained diffusion backbone, facilitating a seamless transition to the second-phase OAT-FM training paradigm. The implementation details are in Appendix C.1.
When deriving the coupling by computing the Wasserstein distance in (3), the FM-based model maintains its low FID score on CIFAR-10 while EDM suffers severe performance degradation no matter what upper-level objective is. For the upper-level objective, we can find that $\mathcal{L}_{\text{OAT}}$ works better than $\mathcal{L}_{\text{CFM}}$ consistently for both FM- and EDM-based models. These results demonstrate that each component of OAT-FM plays an indispensable role in effectively refining flows and enhancing model performance.
Remark. Our OAT-based coupling improves EDM performance, whereas the classic OT-based coupling results in catastrophic performance degradation. This phenomenon reveals the essential difference between OT and OAT in the velocity smoothness. In particular, EDM learns a score function under the assumption that the noise at time $t$ is independent and isotropic Gaussian noise. When reformulating EDM as an FM model (See Appendix C.1), this assumption means that the velocity field is smoothed (song2020score). The OT-based coupling, however, achieves static, non-Markovian transport between noise and data in the sample space, without any smoothness constraint on the velocity. In contrast, the OAT-based coupling matches the “sample-velocity” tuples at $t=0$ with those at $t=1$ in the product space $\mathcal{X}\times\mathcal{V}$, solving the acceleration minimization problem in (4) equivalently. The objective of (4) reveals the isotropic Gaussian prior of the acceleration, which leads to smoothed velocity (benamou2019second). Consequently, our OAT-based coupling follows the smoothness assumption and is suitable for refining EDM.
@@ 4.3 Large-scale Conditional Image Generation
(a) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(b) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(c) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
(d) SiT-XL (Left) v.s. SiT-XL + OAT-FM (Right)
Figure 4: The visual comparison for SiT-XL and SiT-XL + OAT-FM when CFG is 4.0.
To verify the feasibility of OAT-FM as a phase-2 training method in practice, given the state-of-the-art image generator SiT-XL (ma2024sit) trained on the ImageNet $\times256$ benchmark (deng2009imagenet), we apply OAT-FM to continually refine the model on the dataset and test it in conditional image generation tasks. We compare the model refined by OAT-FM with the original SiT-XL and other image generators on different metrics, including FID, spatial FID (sFID), Inception Score (IS) (salimans2016improved), and the precision and recall measuring how well the real and generated data manifolds are overlapped with each other. In the inference phase, we apply an ODE sampler and a SDE one, respectively, to generate images with different Classifier-Free Guidance (CFG) scales.
As detailed in Table 5, for the original SiT-XL model derived by 1,400 training epochs, refining it by OAT-FM with merely five more training epochs (48K batches) leads to consistent improvements in FID, sFID, and IS. The precision and recall remain stable. In Figure 3, we plot the performance of SiT-XL with and without OAT-FM while varying the CFG scale from 1.0 (no guidance) to 4.0. The results confirm that OAT-FM delivers consistent improvements across the entire range of scales. Note that, setting a large CFG scale helps generate high-quality images in general (ma2024sit; peebles2023scalable) (although leading to relatively high FID/sFID scores). Based on this fact, the superiority of OAT-FM is significant when CFG is large (e.g., CFG is 4.0), which further demonstrates the effectiveness of OAT-FM in practice.
Interestingly, when starting from the same noise, in some cases (e.g., the “Bottle” and “Parrot” shown in Figure 4), applying OAT-FM improves image details, maintains the main semantic and spatial content created by SiT-XL, and suppresses hallucination. However, in the other cases (e.g., “Lion” and “Car” shown in Figure 4), applying OAT-FM results in the model producing images that are entirely different from those generated by SiT-XL. This phenomenon indicates that the coupling associated with OAT-FM differs from that in the original model. For flow trajectories whose endpoints obey the data distribution well, OAT-FM straightens them with almost the same endpoints. For flow trajectories whose endpoints are undesired (e.g., in the “Lion” generated by SiT-XL), OAT-FM significantly changes their directions, leading to new endpoints that may better fit the data distribution. More visual results are in Appendix C.4.2.
@@ 5 Conclusion
BigGAN-deep (brock2019large) 6.95 7.36 171.4 0.87 0.28
StyleGAN-XL (sauer2022styleganxl) 2.30 4.02 265.1 0.78 0.53
Mask-GIT (chang2022maskgit) 6.18 - 182.1 - -
ADM-G/U (dhariwal2021diffusion) 3.94 6.14 215.8 0.83 0.53
CDM (ho2021cascaded) 4.88 - 158.7 - -
RIN (jabri2023scalable) 3.42 - 182.0 - -
@@
Limitations and future work. There are natural directions to refine the method. OAT-FM benefits from reasonably accurate endpoint velocities, so Currently, training from scratch via OAT-FM can be fragile due to its dependence on velocity information — early velocity estimates are noisy and can misguide the product-space coupling. As a straightforward solution, warm starts using CFM or self-distillation in the spirit of Shortcut (frans2024one) and consistency models (kim2023consistency) may offer simple remedies before handing off to OAT-FM. However, whether the theory of OAT can be used to develop one-step FM method is still an open problem. From a computational standpoint, mini-batch couplings scale quadratically in batch size. In the future, we plan to explore the dual form of OAT problem and develop a more efficient OAT solver to accelerate OAT-FM further.
@@ B.1 Derivation of (9)
In the context of FM, we have $\pi(z_{0},z_{1})=\pi(x_{0},v_{0},x_{1},v_{1})=\pi_{x}(x_{0},x_{1})\pi(v_{0}|x_{0})\pi(v_{1}|x_{1})$, where $\pi_{x}(x_{0},x_{1})\in\Pi(\rho_{0},\rho_{1})$ is the marginal coupling corresponding to sample pairs and $\pi(\cdot|x_{t})=\delta_{v_{\theta}(x_{t},t)}$ is the Dirac measure determined by the flow model $v_{\theta}$. Based on the decomposition that $\pi(z_{0},z_{1})=\pi_{x}(x_{0},x_{1})\delta_{v_{\theta}(x_{0},0)}(v_{0})\delta_{v_{\theta}(x_{1},1)}(v_{1})$, we can merely optimize $\pi_{x}(x_{0},x_{1})$ and reformulate the lower-level problem in (7) as
@@
The inequality in the above derivation is because we impose the decomposable structure on $\pi$, which shrinks its feasible domain from $\Pi(\mu_{0},\mu_{1})$ to $\Pi^{\prime}(\mu_{0},\mu_{1}):=\{\pi\;|\;\pi(z_{0},z_{1})=\pi_{x}(x_{0},x_{1})\delta_{v_{\theta}(x_{0},0)}(v_{0})\delta_{v_{\theta}(x_{1},1)}(v_{1}),\;\pi_{x}=\iint_{v_{0},v_{1}}\pi,\mbox{ and }\pi\in\Pi\}$.
@@
where $\langle\cdot,\cdot\rangle$ denotes inner product, $\mathbf{T}$ is the coupling matrix, and $\mathbf{C}=[c_{ij}]\in\mathbb{R}^{B\times B}$ is the cost matrix, whose element $c_{ij}=12\|x_{1,i}-x_{0,j}-\frac{v_{0,j}+v_{1,i}}{2}\|_{2}^{2}+\|v_{1,i}-v_{0,j}\|_{2}^{2}$. Following the OT-CFM rationale, the problem can be cast as a linear program with computational complexity $\mathcal{O}(B^{3}\log\|\mathbf{C}\|_{\infty})$. The optimizer can be further approximated by adding an entropic regularizer of $\mathbf{T}$ weighted by $\epsilon$, i.e., $\epsilon\langle\mathbf{T},\log\mathbf{T}\rangle$. This allows the problem to be solved efficiently by Sinkhorn method, whose complexity is $\mathcal{O}(B^{2}\log B)$, and the exact OT result can be recovered by taking $\epsilon$ sufficiently small. We refer to peyre2019computational for further details.
@@ Appendix C Detailed Experiment Settings
Model Paradigm Objective Path / Coupling Architecture Space
EDM (karras2022elucidating) Diffusion Score Matching VP/VE Noise Schedule U-Net Pixel
FM (lipman2023flow) Flow Flow Matching Gaussian-Data Path U-Net Pixel
I-CFM (tong2024improving) Flow Conditional FM Linear Path (Independent) U-Net Pixel
VP-CFM (albergo2023stochastic) Flow Conditional FM Trigonometric Path U-Net Pixel
SB-CFM (tong2024improving) Flow Conditional FM Brownian Bridge Path U-Net Pixel
OT-CFM (tong2024improving) Flow Conditional FM Linear Path (OT Coupling) U-Net Pixel
SiT-XL (ma2024sit) Flow Velocity Matching Linear / GVP Path DiT Latent
@@ C.1 Conversion from VP/VE Models to Flow Matching
Following the work in (lee2024improving), our training process begins with a pre-trained diffusion model. Specifically, we use models trained with the Elucidating the Design Space of Diffusion-Based Generative Models (EDM) framework (karras2022elucidating), which provides a unified perspective on score-based models including Variance Preserving (VP) and Variance Exploding (VE) SDEs. To adapt this model for flow matching, we re-parameterize its input and output to function as a velocity field predictor. This allows us to initialize our model with a strong, pre-trained diffusion backbone, facilitating a seamless transition to the second-phase OAT-FM training paradigm.
EDM Preconditioning. The EDM model takes a noise-corrupted input $x_{\sigma}=x_{1}+\sigma z$ (where $z\sim\mathcal{N}(0,I)$) and a noise level $\sigma$, is trained to denoise it. Here, the noise at time $t$ is assumed to be isotropic Gaussian and independent of the data. Accordingly, the forward process of EDM is Markovian, implying a smooth score function and a smooth velocity field (song2020score). The denoised output $D_{\theta}(x_{\sigma},\sigma)$ is an estimate of the original data $x_{1}$ and is formulated using a preconditioning scheme: